Find angle between the planes r.(i-2j-2k)=1 and r.(3i-6j+2k)=0

Grade:12

1 Answers

Arun
6 years ago
Dear Abhi
 
Cos @ = (n1 . n2)/ |n1| * |n2|
= Cos @.= 11/ 3* 7 = 11/21
 
Regards
Arun (askIITians forum expert)